# A Survey of World Models for Autonomous Driving

### v3

@@ 3.1.4 PC-based Generation
CNN‐based PC Generation. PCP [98] stacks consecutive range images into a 3D spatio-temporal volume and applies 3D CNNs that fuse geometry with motion cues for consistent temporal modelling. 4DOcc [100] discretizes historical scans into 4D occupancy grids and leverages lightweight CNN temporal modules to forecast future occupancy, lowering annotation demands while preserving scene fidelity.
@@
Transformer‐based PC Generation. PCPNet [101] enriches range-image sequences by inserting self-attention layers that capture long-range dependencies, producing temporally consistent point clouds with stronger contextual cues. ViDAR [20] frames vision-LiDAR pre-training as an autoregressive task: it embeds images and actions with a Transformer encoder and decodes future point sequences causally, providing geometry-aware priors that benefit downstream BEV generation. Building on this, HERMES [102] fuses multi-view images and textual prompts into BEV embeddings through causal attention, then employs differentiable voxel rendering, constrained by motion and collision losses, to synthesize high-fidelity point clouds that respect scene dynamics.
@@
Diffusion‐based PC Generation. LiDARGen [103] formulates LiDAR synthesis as a score‐matching diffusion process on equirectangular range images, producing physically plausible and controllable samples. Copilot4D [105] discretizes LiDAR observations with a VQ‐VAE and conducts discrete diffusion to autoregressively sample future sweeps, enabling unsupervised world‐model learning. RangeLDM [106] denoises latent representations of range‐view images to reconstruct accurate point clouds, combining latent diffusion with Hough‐voting projections.
@@
Other PC Generation Methods. Early GAN‐based frameworks [118] support unconditional point‐cloud synthesis, while lidarGeneration [107] specifically mimics real‐world LiDAR statistics via adversarial training on range‐image projections. SPFNet [109] and S2net [110] employ LSTMs on sequential range maps with explicit compensation for sensor calibration. Lidarsim [7] couples physical raycasting with learned noise models to generate near‐realistic scenes. UltraLiDAR [108] uses discrete latent tokens to preserve structural and semantic consistency. Neural‐field methods (i.e., NFL [111] and Nerf‐LiDAR [112]) extend NeRF paradigms to LiDAR, enabling novel‐view synthesis under dynamic and partially observed conditions.
TABLE III: Overview of representative learning-based planners (§3.2.1).
@@
Model-free RL PFBD[13] Electronics’19 TD3 policy net Gap + Speed + Curvature Manoeuvre cmd. Internal[13]
[…]
AdaptiveDriver[119] ICRA’24 BehaviorNet + MPC Agent histories + HD-map Trajectory nuPlan[65]
[…]
AdaWM[123] ICLR’25 Low-rank adaptive WM State + reward Trajectory nuPlan[65]
[…]
Dream2Drive[124] arXiv’25 Diff. physics WM State + goal Trajectory WOMD[125]
[…]
@@
Autoreg. Transformer DrivingGPT[127] arXiv’24 MM GPT-style decoder Image + ego state Action tokens nuPlan[65]
VaViM[128] arXiv’25 Video-Gen. Transformer Video + ego Action tokens OpenDV[53]& nuPlan[65]& nuScenes[36]
[…]
@@
Volume-based Hybrid Cost Volume DSDNet[131] ECCV’20 CNN cost-volume Image + BEV map Trajectory set nuScenes[36]
@@
ST-P3[27] ECCV’22 Spatio-temp. CNN Image seq. Trajectory set nuScenes[36]
@@
MP3[28] CVPR’21 Unified map–predict–plan Multi-modal Trajectory set nuScenes[36]
@@
NEAT[132] ICCV’21 Attention fields Image + HD-map Trajectory set nuScenes[36]
@@
NMP[133] CVPR’19 Neural motion planner LiDAR + HD-map Trajectory set Internal[133]
@@
PPP[115] ECCV’20 Semantic BEV planner BEV semantics Trajectory set nuScenes[36]
[…]
Summary. Table II shows different modern LiDAR generation frameworks. Several observations can be drawn:
[…]
3.2 Behavior Planning for Intelligent Agents
Behavior planning translates a rich, ever-changing scene understanding into a safe, comfortable, goal-directed trajectory. Most methods first sample diverse, kinematically feasible motions and then, guided by semantic predictions and tactical cues, select the most likely candidate [28, 131, 134]. This likelihood-based scoring fuses predicted dynamics, traffic rules, and comfort metrics, yielding a path that adapts smoothly to evolving conditions while meeting stringent safety requirements.
3.2.1 Learning-based Planning
Early autonomous-vehicle planners relied on rule-based heuristics, but these brittle hand-crafted pipelines struggle with the combinatorial variety of urban traffic and become hard to maintain as the rule set expands [135, 136, 137]. Learning-based motion planning gained traction once large-scale driving logs and fast GPUs became available [6]. Modern architectures ingest multimodal sensor streams (i.e., LiDAR, RADAR, GPS signals [138] and camera frames [139, 140]) to output lane-change choices [141, 142], future state distributions [143], reference paths [144], or direct control commands such as steering and throttle [145, 146, 147].
Offline supervision or reinforcement learning allows these networks to keep online inference fast and to adapt to new scenes when supplied with sufficient diverse data. Compared with rule-based planning, data-driven planners handle uncertainty and interaction richness more gracefully, especially when reinforced by graph neural networks, attention mechanisms, or large language models that encode high-level traffic intent [148, 149, 150]. However, three open issues remain: (i) generalisation: performance still degrades in out-of-distribution weather, geography, or traffic densities; (ii) interpretability: opaque latent policies impede validation and debugging [145, 151, 147]; and (iii) formal safety guarantees: current learning pipelines lack provable bounds required for certification [152]. Addressing these gaps is central to scaling learning-based behavior planning from research prototypes to trustworthy real-world deployment.
RL & MPC Planners. Model-free reinforcement-learning (RL) planner [13] employs a twin-delayed deep deterministic policy gradient (TD3) that maps high-level planning features to manoeuvre commands, delegating the final continuous trajectory to a downstream generator. Model-based RL planners span a spectrum of world-model granularity and safety mechanisms. Early deterministic offline ensembles [34] learn compact latent dynamics and couple them with Model Predictive Control (MPC) to deliver fast, sample-efficient roll-outs and transparent constraint handling, yet they over-commit when aleatoric noise appears in multi-agent traffic. Building on this paradigm, the follow-up work [119] performs planning with adaptive world models, learning a closed-loop social world model and embedding it within MPC – an explicit instantiation of model-based RL tailored to urban traffic. Uncertainty-aware variants [120, 123] inject stochastic ensembles or adaptive low-rank updates, exposing epistemic and aleatoric risk for more robust out-of-distribution reasoning, though at the price of heavier sampling and fine-tuning overheads. Agent-centric and analytic approaches [41, 124] model each actor separately – using attention-based latent intents or differentiable physics – to capture nuanced social interactions, but they rely on accurate tracks or hand-coded dynamics. Latent-space accelerators [126] plan entirely in a low-dimensional latent, enabling orders-of-magnitude faster roll-outs on modest hardware, yet abstract states can omit critical geometry such as traffic-light phase and undermine fine manoeuvres.
LLM-based Planners. Autoregressive multimodal transformers [127, 128] treat driving as next-token prediction over interleaved image-and-action tokens, coupling perception, world modelling and control, and achieving strong nuPlan and NAVSIM results, but they demand high-end multi-GPU compute and still struggle with long-horizon consistency and safety certification. WoTE adds a lightweight bird’s-eye-view predictor that scores sampled trajectories online, boosting safety with minimal latency, yet its fidelity is capped by BEV grid resolution and the quality of the trajectory sampler [129]. SSR compresses dense BEV features into just 16 navigation-guided tokens, cutting FLOPs while maintaining accuracy, though aggressive sparsification risks missing rare but critical actors or occlusions [130]. DriveSim explores GPT-4V-style models as internal simulators for language-conditioned planning, showing promising affordance reasoning but revealing frame-order errors and inconsistent imagined trajectories that limit closed-loop reliability [66].
Volume-based Planners. Hybrid cost-volume methods first sample or generate a diverse set of candidate trajectories and then rank them with a composite cost that fuses hand-crafted comfort or rule-compliance terms with likelihoods predicted from learned occupancy or segmentation maps. To preserve interpretability, such planners explicitly build a cost map, pair it with a trajectory sampler, and execute the candidate with the lowest overall cost. Concretely, cost-volume planners [133, 115, 28, 153] compute a confidence score for each sampled trajectory from the cost volume, guiding the choice toward paths that best balance safety, efficiency, and rule compliance. This approach yields interpretable, safety-oriented plans, but its effectiveness diminishes when the sampler cannot cover the dense combinatorial space of urban traffic [131, 27, 28, 132, 133, 115]. By contrast, occupancy volume method replaces heuristic costs with a learnable head applied to 4D occupancy forecasts, providing richer collision cues and enabling end-to-end fine-tuning, albeit at the expense of higher memory footprints and a reliance on dense occupancy supervision [88].

### v4

@@ 3.1.4. PC-based Generation
CNN‐based PC Generation. PCP (Mersch et al., 2022) stacks consecutive range images into a 3D spatio-temporal volume and applies 3D CNNs that fuse geometry with motion cues for consistent temporal modelling. 4DOcc (Khurana et al., 2023) discretizes historical scans into 4D occupancy grids and leverages lightweight CNN temporal modules to forecast future occupancy, lowering annotation demands while preserving scene fidelity.
@@
Transformer‐based PC Generation. PCPNet (Luo et al., 2023) enriches range-image sequences by inserting self-attention layers that capture long-range dependencies, producing temporally consistent point clouds with stronger contextual cues. ViDAR (Yang et al., 2024a) frames vision-LiDAR pre-training as an autoregressive task: it embeds images and actions with a Transformer encoder and decodes future point sequences causally, providing geometry-aware priors that benefit downstream BEV generation. Building on this, HERMES (Zhou et al., 2025) fuses multi-view images and textual prompts into BEV embeddings through causal attention, then employs differentiable voxel rendering, constrained by motion and collision losses, to synthesize high-fidelity point clouds that respect scene dynamics.
@@
Diffusion‐based PC Generation. LiDARGen (Zyrianov et al., 2022) formulates LiDAR synthesis as a score‐matching diffusion process on equirectangular range images, producing physically plausible and controllable samples. Copilot4D (Zhang et al., 2024b) discretizes LiDAR observations with a VQ‐VAE and conducts discrete diffusion to autoregressively sample future sweeps, enabling unsupervised world‐model learning. RangeLDM (Hu et al., 2024b) denoises latent representations of range‐view images to reconstruct accurate point clouds, combining latent diffusion with Hough‐voting projections. LiDARCrafter (Liang et al., 2025a) extends LiDAR generation to dynamic 4D world modeling by parsing free-form text into an ego-centric scene graph, conditioning a tri-branch diffusion network for layout synthesis, and enforcing temporal coherence via an autoregressive warping module with ego/agent motion priors.
@@
Other PC Generation Methods. Early GAN‐based frameworks (Sallab et al., 2019) support unconditional point‐cloud synthesis, while lidarGeneration (Caccia et al., 2019) specifically mimics real‐world LiDAR statistics via adversarial training on range‐image projections. SPFNet (Weng et al., 2021) and S2net (Weng et al., 2022) employ LSTMs on sequential range maps with explicit compensation for sensor calibration. Lidarsim (Manivasagam et al., 2020) couples physical raycasting with learned noise models to generate near‐realistic scenes. UltraLiDAR (Xiong et al., 2023) uses discrete latent tokens to preserve structural and semantic consistency. Neural‐field methods (i.e., NFL (Huang et al., 2023a) and Nerf‐LiDAR (Zhang et al., 2024d)) extend NeRF paradigms to LiDAR, enabling novel‐view synthesis under dynamic and partially observed conditions.
Summary. Table 2 shows different modern LiDAR generation frameworks. Several observations can be drawn:
3.2. Behavior Planning for Intelligent Agents
Behavior planning translates a rich, ever-changing scene understanding into a safe, comfortable, goal-directed trajectory. Most methods first sample diverse, kinematically feasible motions and then, guided by semantic predictions and tactical cues, select the most likely candidate (Casas et al., 2021; Zeng et al., 2020; Hu et al., 2021a). This likelihood-based scoring fuses predicted dynamics, traffic rules, and comfort metrics, yielding a path that adapts smoothly to evolving conditions while meeting stringent safety requirements.
3.2.1. Learning-based Planning
Early autonomous-vehicle planners relied on rule-based heuristics, but these brittle hand-crafted pipelines struggle with the combinatorial variety of urban traffic and become hard to maintain as the rule set expands (Karnchanachari et al., 2024; Paden et al., 2016). Learning-based motion planning gained traction once large-scale driving logs and fast GPUs became available (Schwarting et al., 2018). Modern architectures ingest multimodal sensor streams (i.e., LiDAR, RADAR, GPS signals (Caltagirone et al., 2017) and camera frames (Bojarski et al., 2016; Richter and Roy, 2017)) to output lane-change choices (Vallon et al., 2017; Huy et al., 2013), future state distributions (Mozaffari et al., 2020), reference paths (Lefevre et al., 2015), or direct control commands such as steering and throttle (Pomerleau, 1989; Lu et al., 2019; Prakash et al., 2021).
Offline supervision or reinforcement learning allows these networks to keep online inference fast and to adapt to new scenes when supplied with sufficient diverse data. Compared with rule-based planning, data-driven planners handle uncertainty and interaction richness more gracefully, especially when reinforced by graph neural networks, attention mechanisms, or large language models that encode high-level traffic intent (Jia et al., 2023a, 2021, b). However, three open issues remain: (i) generalisation: performance still degrades in out-of-distribution weather, geography, or traffic densities; (ii) interpretability: opaque latent policies impede validation and debugging (Pomerleau, 1989; Codevilla et al., 2019; Prakash et al., 2021); and (iii) formal safety guarantees: current learning pipelines lack provable bounds required for certification (Lu et al., 2024). Addressing these gaps is central to scaling learning-based behavior planning from research prototypes to trustworthy real-world deployment.
RL & MPC Planners. Model-free reinforcement-learning (RL) planner (Qian et al., 2019) employs a twin-delayed deep deterministic policy gradient (TD3) that maps high-level planning features to manoeuvre commands, delegating the final continuous trajectory to a downstream generator. AdaptiveDriver (Vasudevan et al., 2024) performs planning with adaptive world models by learning a closed-loop social world model and embedding it into an MPC planner. In contrast, AdaWM (Wang et al., 2025d) formulates adaptive world-model RL: under distribution shift it diagnoses world-model/policy mismatch via a divergence measure and triggers alignment-driven finetuning while execution is through a learned RL policy. Uncertainty-aware variants (Diehl et al., 2023) inject stochastic ensembles or adaptive low-rank updates, exposing epistemic and aleatoric risk for more robust out-of-distribution reasoning, though at the price of heavier sampling and fine-tuning overheads. Agent-centric and analytic approaches (Nachkov et al., 2025) model each actor separately – using attention-based latent intents or differentiable physics – to capture nuanced social interactions, but they rely on accurate tracks or hand-coded dynamics. Latent-space accelerators (Li et al., 2024b) plan entirely in a low-dimensional latent, enabling orders-of-magnitude faster roll-outs on modest hardware, yet abstract states can omit critical geometry such as traffic-light phase and undermine fine manoeuvres. Model-based RL learns a predictive world model and optimizes the policy via imagination roll-outs, improving sample efficiency and closed-loop reliability; recent exemplars include Raw2Drive’s dual-stream alignment between a privileged and a raw-sensor world model (Yang et al., 2025b) and PIWM’s DreamerV3-style individual world model (Gao et al., 2024d) that encodes inter-agent intentions via trajectory prediction, enabling imagination-based RL. Moreover, there has been a growing trend of integrating RL with vision language model (VLM) and vision-language-action (VLA) models in autonomous driving. VL‑SAFE employs offline, world‑model RL guided by a VLM: it labels logs with VLM safety scores, imagines safety‑aware roll‑outs, and trains an actor-critic policy (Qu et al., 2025). IRL‑VLA constructs a reward world model via inverse RL to replace costly simulator-based rewards, and subsequently uses the RWM to guide PPO-based RL training of a VLA policy (Jiang et al., 2025a).
LLM-based Planners. Autoregressive multimodal transformers (Chen et al., 2024; Bartoccioni et al., 2025) treat driving as next-token prediction over interleaved image-and-action tokens, coupling perception, world modelling and control, and achieving competitive nuPlan results, but they demand high-end multi-GPU compute and still struggle with long-horizon consistency and safety certification. WoTE adds a lightweight bird’s-eye-view predictor that scores sampled trajectories online, boosting safety with minimal latency, yet its fidelity is capped by BEV grid resolution and the quality of the trajectory sampler (Li et al., 2025d). SSR compresses dense BEV features into just 16 navigation-guided tokens, cutting FLOPs while maintaining accuracy, though aggressive sparsification risks missing rare but critical actors or occlusions (Li and Cui, 2025). DriveSim explores GPT-4V-style models as internal simulators for language-conditioned planning, showing promising affordance reasoning but revealing frame-order errors and inconsistent imagined trajectories that limit closed-loop reliability (Sreeram et al., 2024). FSDrive (Zeng et al., 2025) reframes intermediate reasoning from text CoT to a spatio-temporal visual CoT: a VLM first acts as a world model to imagine a unified future frame overlaying future lanes and 3D boxes (spatial) together with an ordinary future frame (temporal), and then, conditioned on these visual CoT predictions, outputs the trajectory as an inverse-dynamics step; a unified pretraining scheme activates visual generation while preserving understanding.
Volume-based Planners. Hybrid cost-volume methods first sample or generate a diverse set of candidate trajectories and then rank them with a composite cost that fuses hand-crafted comfort or rule-compliance terms with likelihoods predicted from learned occupancy or segmentation maps. To preserve interpretability, such planners explicitly build a cost map, pair it with a trajectory sampler, and execute the candidate with the lowest overall cost. Concretely, cost-volume planners (Zeng et al., 2019; Sadat et al., 2020; Casas et al., 2021; Phan-Minh et al., 2020) compute a confidence score for each sampled trajectory from the cost volume, guiding the choice toward paths that best balance safety, efficiency, and rule compliance. This approach yields interpretable, safety-oriented plans, but its effectiveness diminishes when the sampler cannot cover the dense combinatorial space of urban traffic (Zeng et al., 2020; Hu et al., 2022a; Casas et al., 2021; Chitta et al., 2021; Zeng et al., 2019; Sadat et al., 2020). By contrast, occupancy volume method replaces heuristic costs with a learnable head applied to 4D occupancy forecasts, providing richer collision cues and enabling end-to-end fine-tuning, albeit at the expense of higher memory footprints and a reliance on dense occupancy supervision (Yang et al., 2025c).
Other Planners. MILE (Hu et al., 2022b) introduces a model-based imitation learning (IL) framework that leverages 3D geometry as an inductive bias to jointly learn static scenes, surrounding agents, and ego-vehicle states from high-resolution camera data. It acquires a generative world model and driving policy from offline expert logs, and employs action-conditioned rollouts to simulate diverse future scenarios. BEV-Planner (Li et al., 2024c), a label-free baseline, directly regresses trajectories through cross-attention between ego-vehicle queries and temporal BEV features, with optional integration of ego status. UncAD (Yang et al., 2025d) makes the BEV map itself probabilistic: a Map Uncertainty Estimation (MUE) head predicts BE BEVV map points and their Laplace parameters, then an Uncertainty-Guided Prediction & Planning (UGPnP) stage generates multi-modal trajectories, and an Uncertainty–Collision-Aware Selection (UCAS) ranks them by collision risk and map confidence.
Table 3. Overview of representative learning-based planners (§3.2.1). Please note that methods that use VLM/VLA only to assist RL training or safety scoring are listed under RL & MPC planners; methods where a (V)LLM acts as the planner at inference appear under LLM-based planners.
@@ 3.2.1. Learning-based Planning
Model-free RL PFBD(Qian et al., 2019) Electronics’19 TD3 policy net Gap + Speed + Curvature Manoeuvre cmd. Internal(Qian et al., 2019)
[1pt/1pt] Adaptive World Model AdaptiveDriver(Vasudevan et al., 2024) arXiv’24 BehaviorNet + MPC Agent histories + HD-map Trajectory nuPlan(Caesar et al., 2021)
AdaWM(Wang et al., 2025d) ICLR’25 Low-rank adaptive WM State + reward Trajectory nuPlan(Caesar et al., 2021)
[1pt/1pt] Uncertainty-aware UMBRELLA(Diehl et al., 2023) RA-L’23 Ensemble + Unc. propag. Vehicle state history Trajectory NGSIM(Henaff et al., 2018) + CARLA(CARLA Team, 2022)
[1pt/1pt] Agent-centric/Analytic Dream2Drive(Nachkov et al., 2025) arXiv’25 Diff. physics WM State + goal Trajectory WOMD(Ettinger et al., 2021)
[1pt/1pt] Latent accelerator Think2Drive(Li et al., 2024b) ECCV’24 Latent-space planner Latent WM Trajectory CARLA-V2(CARLA Team, 2022)
[1pt/1pt] Model-base RL Raw2Drive(Yang et al., 2025b) arXiv’25 Neural planner + Policy RL Multi-view cameras + Ego state Low-level actions CARLA V2(CARLA Team, 2022) + Bench2Drive(Jia et al., 2024)
PIWM(Gao et al., 2024d) TIV’25 Actor–critic RL Agent histories/relations + Ego Action + Trajectory I-SIM(Zhang et al., 2022)
[1pt/1pt] VLM/VLA VL-SAFE(Qu et al., 2025) arXiv’25 VLM + WM + offline safe RL Image + Text + Offline expert logs Low-level actions/Policy Internal(Qu et al., 2025)
IRL-VLA(Jiang et al., 2025a) arXiv’25 IRL + PPO RL + VLA policy Multi-view images/video + Ego state Trajectory + Actions NAVSIM v2(Dauner et al., 2024)
@@
Autoreg. Transformer DrivingGPT(Chen et al., 2024) arXiv’24 MM GPT-style decoder Image + ego state Action tokens nuPlan(Caesar et al., 2021) & NAVSIM(Dauner et al., 2024)
VaViM(Bartoccioni et al., 2025) arXiv’25 Video-Gen. Transformer Video + ego Action tokens OpenDV(Yang et al., 2024b)& nuPlan(Caesar et al., 2021)& nuScenes(Caesar et al., 2020)
[1pt/1pt] BEV scorer WoTE(Li et al., 2025d) arXiv’25 BEV head + sampler BEV grid + cand. traj. Score vector NAVSIM(Dauner et al., 2024)&Bench2Drive(Jia et al., 2024)
[1pt/1pt] Sparse-token WM SSR(Li and Cui, 2025) ICLR’25 Sparse token WM BEV features Action tokens nuScenes(Caesar et al., 2020) & CARLA(Dosovitskiy et al., 2017a)
[1pt/1pt] Vision-LLM sim. DriveSim(Sreeram et al., 2024) arXiv’24 GPT-4V probe Lang. cmd. + frames Sim. trajectory Internal(Sreeram et al., 2024)
[1pt/1pt] Visual CoT FSDrive(Zeng et al., 2025) arXiv’25 VLM + visual CoT Surround cameras + Nav/ego Trajectory + future frames nuScenes(Caesar et al., 2020)
@@
Volume-based Hybrid Cost Volume DSDNet(Zeng et al., 2020) ECCV’20 CNN cost-volume Image + BEV map Trajectory set nuScenes(Caesar et al., 2020) & ATG4D(Zeng et al., 2020) & Precog(Rhinehart et al., 2019)
@@
ST-P3(Hu et al., 2022a) ECCV’22 Spatio-temp. CNN Image seq. Trajectory set nuScenes(Caesar et al., 2020) & CARLA(Dosovitskiy et al., 2017a)
@@
MP3(Casas et al., 2021) CVPR’21 Unified map–predict–plan Multi-modal Trajectory set nuScenes(Caesar et al., 2020)
@@
NEAT(Chitta et al., 2021) ICCV’21 Attention fields Image + HD-map Trajectory set nuScenes(Caesar et al., 2020)
@@
NMP(Zeng et al., 2019) CVPR’19 Neural motion planner LiDAR + HD-map Trajectory set Internal(Zeng et al., 2019)
@@
PPP(Sadat et al., 2020) ECCV’20 Semantic BEV planner BEV semantics Trajectory set nuScenes(Caesar et al., 2020)
[1pt/1pt] Occupancy volume Occ-WM(Yang et al., 2025c) AAAI’25 4D occ. WM + $Q$-net 4D occupancy Trajectory nuScenes(Caesar et al., 2020) +  Occ3D(Tian et al., 2024)
Others
Imitation Learning MILE(Hu et al., 2022b) NIPS’22 Latent dynamics + Policy Image + ego actions Action + BEV Segmentation CARLA expert logs(Dosovitskiy et al., 2017a)
Ego Status BEV-Planner(Li et al., 2024c) CVPR’24 Ego-query cross-attn + MLP Images + Ego status Trajectory nuScenes(Caesar et al., 2020)
Uncertainty-aware UncAD(Yang et al., 2025d) arXiv’25 MUE (Laplace BEV) + UGPnP + UCAS Multi-view images + Ego state Trajectory nuScenes(Caesar et al., 2020)
Summary. Table 3 shows compares major learning-based planning paradigms. Several observations can be drawn:
3.2.2. Rule-based Planning
Rule-based planners endure because they guarantee safety and are easy to audit (González et al., 2015; Zhou et al., 2022). They treat motion as a deterministic optimisation over explicit dynamics and hand-crafted costs, yielding transparent behaviour. Four core families dominate: car-following laws (e.g., IDM), sampling planners (RRT, state lattice), model-predictive control, and potential-field navigation. Modern stacks combine these blocks for interpretability and safety, yet modular fusion can lose global optimality and fixed heuristics struggle in dense, fast-changing traffic.
Car-following Models. Physics-inspired longitudinal models translate the instantaneous gap, speed and closing rate into throttle or brake commands. The Intelligent Driver Model and its calibrated siblings capture free-flow, synchronized and stop‐and-go regimes with a single continuous formula (Treiber et al., 2000). Recent work samples multiple IDM roll-outs and selects the minimum-cost, safety-compliant candidate to improve comfort without losing closed-form guarantees (Dauner et al., 2023). Because these laws are purely algebraic, production-grade Adaptive Cruise Control (ACC) systems can execute them at tens-to-hundreds Hz frequencies. However, their inherently short prediction horizon is a key limitation, so modern planners augment analytic commands with map priors or predictive MPC to anticipate intersections and cut-in maneuvers better.
Sampling-based Planners. Sampling-based planners probe the vehicle’s configuration space with a sequence of samples, accept only collision-free continuations, and then rank feasible motion sequences with handcrafted costs. They split into two sub-families that trade optimality for speed in different ways. Random sampling methods grow a tree or roadmap toward stochastically chosen states, guaranteeing probabilistic completeness and – when augmented with rewiring – asymptotic optimality. Classic Rapidly-Exploring Random Trees (RRT) (LaValle, 1998; Kuffner and LaValle, 2000) realise millisecond-level feasibility checks for single-query problems. Their optimal variants, RRT and RRG, refine paths online by incremental rewiring (Karaman and Frazzoli, 2010, 2011). Closed-loop extensions embed vehicle dynamics for high-speed manoeuvres, as demonstrated on MIT’s “Talos” car in the DARPA Urban Challenge (Kuwata et al., 2009; Buehler et al., 2009) and in multi-agent settings (Pongpunwattana and Rysdyk, 2004). Deterministic sampling methods pre-tile the state–time space with kinematically feasible motion primitives, then apply graph search to pick the lowest-cost, collision-free sequence. Spatiotemporal lattices and their conformal highway variant capture curvature and speed limits explicitly (Ziegler and Stiller, 2009; McNaughton et al., 2011). Polynomial libraries (cubic, quintic, B-spline) further reduce run-time by storing analytically smooth segments that satisfy dynamic bounds (Li et al., 2015; Chu et al., 2012). Discretised terminal manifolds shrink search depth for time-critical street scenarios (Werling et al., 2012), and the Stanford “Junior” entry shows how such libraries integrate into a full urban-driving stack (Montemerlo et al., 2008). Because the candidate set is finite, deterministic planners sacrifice global optimality for predictable timing, making them popular for production autonomous vehicle stacks.
Continuous Optimisation. Continuous optimisation planners rebuild vehicle paths in continuous space each control cycle, casting curvature and speed as variables in a quadratic or nonlinear program constrained by dynamics and obstacle corridors. Warm-starting with the previous solution yields smooth, near-optimal trajectories without discrete motion libraries. Apollo’s Expectation–Maximisation planner alternates lateral smoothing with longitudinal speed optimisation, whereas Rastelli’s continuous-curvature scheme minimises jerk and curvature variation along Bézier splines. Both show that convex or near-convex formulations can meet real-time demands while overcoming the sub-optimality and roughness of sampling-based schemes, though at the cost of continual optimisation (Fan et al., 2018; Pérez Rastelli et al., 2014).
Artificial Potential-field Navigation. Attractive and repulsive potentials characterise the goal as an energy well and obstacles as peaks, letting a robot follow the negative gradient in real time (Khatib and Le Maitre, 1978; Khatib, 1986). Newton-style updates hasten escape from saddle points (Ren et al., 2006), and adaptive gain tuning lessens oscillations on public roads (Bounini et al., 2017). Blending the field with sigmoid curves limits curvature spikes in dense traffic (Lu et al., 2020), while embedding the gradient in a torque-based proportional–derivative steering loop unifies motion planning with low-level control (Galceran et al., 2015). Although local minima persist, coupling potential-field guidance with higher-level sampling or MPC preserves rapid collision avoidance without forfeiting global optimality (González et al., 2015; Zhou et al., 2022).
Summary. Table 4 compares four major rule-based motion-planning paradigms. Several observations follow: